OpenStudy (anonymous):

Which statement is true about the difference square root of 7 - square root of 28? It is rational and equal to -2. It is rational and equal to 0. It is irrational and equal to -2square root of 7. It is irrational and equal to square root of 7.

OpenStudy (anonymous):

@jinxhead20 , @e.mccormick , @magepker728 , @mathmale , @Mokeira , @Burningitdown

OpenStudy (one098):

\[\sqrt{7}-\sqrt{28}\] What are the square roots of each?

OpenStudy (anonymous):

0 and 14?

OpenStudy (mokeira):

\[\sqrt{28}=\sqrt{7\times4}=2\sqrt{7}\] therefore \[\sqrt{7}-2\sqrt{7}=?\]

OpenStudy (mokeira):

imagine it this way b-2b wher b is sqrt 7

OpenStudy (anonymous):

wouldn't that be rational and equal to 0 ?

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

it is equal to \[-\sqrt{7}\] but i do not see this in choices